OpenStudy (anonymous):

Simplify:square root of 5 open parentheses 8 plus 3 square root of 6 close parentheses 8 square root of 5 end root plus 3 square root of 30 5 square root of 10 130 None of the above

OpenStudy (anonymous):

Distribute. 8√5 + 3√5 x 6 = 8√5 + 3√30

jhonyy9 (jhonyy9):

sqrt5 (8+3sqrt6) = 8sqrt5 + 3sqrt5 *sqrt6 = 8sqrt5 +3sqrt(5*6) = 8sqrt5 +3sqrt30